The standard cost card for one unit of a finished product shows the following:

Standard Quantity or Hours Standard Price or Rate
Direct materials 12 feet $ ? per foot
Direct labor 1.5 hours $ 12 per hour
Variable manufacturing overhead 1.5 hours $ 8 per hour

If the total standard variable cost for one unit of finished product is $78, then the standard price per foot for direct materials is:

Multiple Choice

  • $2

  • $3

  • $4

  • $5
Answer #1

Direct labor cost = 1.5*12 = $18 per unit

Variable manufacturing overhead = 1.5*8 = $12 per unit

Direct material cost = $78-18-12 = $48 per unit

Standard price per foot for direct material = $48/12 = $4 per foot